OpenStudy (anonymous):

Find the product of (x + 3)(x − 3). x2 − 6x + 9 x2 + 6x + 9 x2 + 9 x2 − 9

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (mayankdevnani):

\[\huge \bf (x+3)(x-3)\] use \[\huge \bf (a-b)(a+b) ~identity\] and this identity states that:- \[\huge \bf (a-b)(a+b)=a^2-b^2\] where a=3 and b=3

OpenStudy (mayankdevnani):

can you solve it? @supanick14

OpenStudy (skullpatrol):

guide the asker to a solution

OpenStudy (cggurumanjunath):

d is answer(a^2-b^2) hence (x^2-9)

OpenStudy (skullpatrol):

guide the asker to a solution

OpenStudy (mayankdevnani):

@skullpatrol i am guiding.I am not answering directly.

OpenStudy (anoop27):